# Robust Scene Coordinate Regression via Geometrically-Consistent Global Descriptors

### v1

@@ Abstract
Figure 1: Performance overview on Aachen Day/Night dataset. Our method achieves significant improvements over existing learning-based approaches while maintaining comparable memory efficiency. Compared to R-Score [rscore], we achieve $6.1\%$ higher accuracy on night-time images at the $0.25\text{m}$ / $2^{\circ}$ threshold and 2.2% average improvement across all evaluation thresholds (detailed results in Table 1). This performance narrows the gap between learning-based and traditional structure-based methods while preserving the memory advantages of coordinate regression approaches.
@@ 1 Introduction
Visual localization, the task of determining a camera’s position and orientation from images alone, represents a core challenge in computer vision with critical applications in augmented reality, autonomous navigation, and robotics. While traditional structure-based approaches achieve robust performance through extensive feature matching and geometric verification, they impose significant computational and memory burdens that limit their practical deployment [hloc, active_search].
Scene coordinate regression (SCR) methods [brachmann2023accelerated, GLACE2024CVPR, brachmann2018learning, focustune, DBLP:conf/cvpr/ShottonGZICF13, DBLP:conf/cvpr/ValentinNSFIT15] have emerged as an attractive alternative, excelling in small-scale environments in both accuracy and memory efficiency. However, they face perceptual aliasing as a fundamental challenge in larger environments. In such environments, visually similar landmarks, such as repeated architectural elements, often generate nearly identical local descriptors. This similarity creates ambiguity that makes it difficult to associate descriptors with their correct 3D locations and severely hampers the training effectiveness.
To address the challenge of scene ambiguity, recent works [GLACE2024CVPR, rscore] have leveraged global context to disambiguate similar local features by incorporating global descriptors. R-Score [rscore] extracts geometric relationships from covisibility graphs via the node2vec optimizer [grover2016node2vec]. In these graphs, edges represent spatial proximity between images, with images in close proximity in the covisibility graph sharing similar embeddings. These embeddings, when concatenated with local descriptors, enable effective disambiguation across scenes and substantially boost the performance of learning-based SCR methods.
However, R-Score derives global descriptors from geometric cues alone, decoupling them from visual content. This separation creates two fundamental problems: First, rich visual information that could enhance descriptor discriminability is ignored. Second, the system is vulnerable to the noise in the underlying covisibility graph, which is computed from potentially unreliable overlap scores [rau2020predicting]. When incorrect connections between dissimilar image pairs are formed, the resulting graph embeddings become partially misleading, degrading both descriptor quality and training stability.
@@
Our aggregator learns these geometrically-consistent global descriptors through a training strategy that combines an efficient batch mining strategy relying solely on the overlap scores, and thus eliminating the need for external place labels, with a modified Generalized Contrastive Loss [leyva2023gcl] that uses a smoother distance function.
Experimental validation on challenging benchmarks demonstrates the effectiveness of our approach. As shown in Figure 1, our method achieves a $6.1\%$ improvement at the $0.25\text{m}$ / $2^{\circ}$ threshold compared to R-Score [rscore], with only a 6 MB increase in memory. This gain further narrows the gap to traditional structure-based methods while preserving the memory advantages that make SCR models attractive for practical deployment.
@@
    We demonstrate consistent improvements across challenging benchmarks: compared to R-Score [rscore], our method achieves average accuracy gains of $2.2\%$ and $2.5\%$ on the Aachen Day/Night [sattler2018benchmarking] and the Hyundai Department Store datasets [naverdatasets], respectively, while maintaining similar computational and memory efficiency.
@@ 2.1 Structure-based visual localization
Structure-based methods establish visual localization by matching query images against pre-built 3D scene representations. Early approaches [DBLP:conf/eccv/LiSH10, active_search, sattler2011fast] directly match 2D features with 3D point clouds to establish 2D-3D correspondences. These approaches typically construct a descriptor codebook for each 3D point by averaging descriptors from all database pixels in which the point appears. Direct matching methods can benefit from sophisticated search strategies based on both 2D-to-3D and 3D-to-2D search for additional matches [active_search, sattler2011fast]. However, they remain vulnerable to perceptual aliasing, where visually similar but spatially distinct 3D points produce incorrect correspondences.
Hierarchical approaches [peng2021megloc, hloc] address this limitation by leveraging global descriptors to streamline and refine the matching process. These approaches first retrieve database images similar to the query image using global descriptors, and then establish 2D-2D feature correspondences between retrieved and query images to obtain 2D-3D correspondences. This two-stage process significantly improves accuracy compared to direct matching.
Despite their strong performance, both direct and hierarchical approaches impose substantial memory requirements. They must maintain either complete database image descriptors [hloc, peng2021megloc] or the descriptors for 3D points [DBLP:conf/eccv/LiSH10, active_search, sattler2011fast]. Although recent advances [wang2024mad, laskar2024differentiable] have explored descriptor compression, memory requirements of structure-based algorithms still scale linearly with environment size, unlike learning-based counterparts. This motivates our focus on learning-based alternatives that offer more favorable memory characteristics.
@@ 2.2 Learning-based visual localization
Scene coordinate regression (SCR) models represent a fundamentally different approach to visual localization, directly predicting 3D scene coordinates for pixels in query images [DBLP:conf/cvpr/ShottonGZICF13, DBLP:conf/cvpr/ValentinNSFIT15, DBLP:journals/pami/BrachmannR22, cavallari2017fly, li2020hierarchical]. Early SCR approaches employed random forests [DBLP:conf/cvpr/ShottonGZICF13, DBLP:conf/cvpr/ValentinNSFIT15], while subsequent approaches have adopted convolutional neural networks [DBLP:journals/pami/BrachmannR22, brachmann2016uncertainty, brachmann2018learning, brachmann2019expert, do2022learning, huang2021vs] and fully-connected architectures [brachmann2023accelerated, focustune] to improve the prediction accuracy and computational efficiency.
However, SCR approaches have historically underperformed in large-scale environments due to perceptual aliasing and ambiguous visual patterns. This fundamental challenge has motivated recent efforts to incorporate global context into SCR architectures. Recent works [GLACE2024CVPR, rscore] integrate global descriptors to provide contextual information for disambiguation. GLACE [GLACE2024CVPR] incorporates off-the-shelf descriptors with noise augmentation during training, but achieves limited robustness in large scenes due to insufficient consistency between descriptors of the same location. R-Score [rscore] improves scalability by learning global descriptors from a covisibility graph using node2vec embeddings [grover2016node2vec]. While R-Score demonstrated significant improvements, its graph-based representation operates independently of visual content in the image space, potentially missing important visual cues that could enhance descriptor discriminability.
@@ 2.3 Local descriptors
Local feature descriptors form the foundation of many visual localization systems, enabling the identification and description of consistent pixels under variations in lighting, viewpoint, and scale. Classical methods [DBLP:journals/ijcv/Lowe04, DBLP:conf/eccv/BayTG06] detect invariant keypoints that can be reliably matched across different views. These approaches achieve strong real-world performance with computational efficiency, making them popular choices for structure-based localization systems [DBLP:conf/eccv/LiSH10, active_search, DBLP:conf/cvpr/IrscharaZFB09, sattler2016efficient].
More recent work has shifted toward learning-based local descriptors [detone2018superpoint, noh2017large, dusmanu2019d2, edstedt2024dedode]. Noh et al. [noh2017large] introduces an attention mechanism to highlight semantically meaningful local features while also estimating their confidence. SuperPoint [detone2018superpoint] learns keypoint detection and description on a synthetic dataset containing simple geometric shapes. D2-Net [dusmanu2019d2] employs a unified convolutional neural network for both dense feature description and keypoint detection, postponing the detection process to produce more stable keypoints compared to traditional methods that rely on early detection of low-level image structures. R2D2 [revaud2019r2d2] jointly learns keypoint detection, descriptor extraction, and discriminativeness prediction. This integrated approach reduces the impact of ambiguous regions, leading to more robust and reliable keypoints and descriptors. DeDoDe [edstedt2024dedode] trains a detector using tracks from large-scale structure-from-motion and learns descriptors by optimizing a mutual nearest neighbor objective over keypoints. DeDoDe achieves strong performance in feature matching and has proven particularly effective for scene coordinate regression models [rscore].
@@ 2.4 Global descriptors
Global descriptors enable finding the most similar database images to an input query image by encoding image-wide visual characteristics into compact representations [masone2021survey, DBLP:journals/trob/LowryS0LCCM16]. Current systems often reduce this problem to a similarity search in a $d$-dimensional descriptor space [hausler2021patch, DBLP:conf/cvpr/ArandjelovicGTP16, ali2023mixvpr, berton2023eigenplaces, DBLP:conf/mir/MohedanoMOSMN16, tolias2013aggregate, Izquierdo_CVPR_2024_SALAD, izquierdo2024close, leyva2023gcl].
Global descriptors can be obtained by aggregating either local descriptors [DBLP:conf/mir/MohedanoMOSMN16, tolias2013aggregate], multiple convolutional neural network layers [DBLP:conf/cvpr/ArandjelovicGTP16, ali2023mixvpr, berton2023eigenplaces], or DINO features [dino, dinov2] via optimal transport [Izquierdo_CVPR_2024_SALAD], into a single global descriptor vector. In the context of scene coordinate regression, global descriptors have emerged as crucial components providing contextual information for disambiguating local descriptors in the presence of perceptual aliasing [GLACE2024CVPR, rscore]. Our work advances this trend by introducing a dedicated aggregator module that produces high-quality global descriptors through joint optimization with both geometric and visual constraints, leading to further improvements in scene coordinate regression performance.
@@ 2.5 Foundation models
Foundation models [wang2024dust3r, wang2025vggt, mast3r, reloc3r] have demonstrated strong capabilities in estimating various 3D properties, such as point clouds, depth maps, and camera poses, from just a few images, thanks to their high-quality representations. Dust3r [wang2024dust3r] pioneered this line of work by introducing pointmap regression using powerful pretrained features [weinzaepfel2023croco]. Building on this trend, Wang et al. [wang2025vggt] proposed the Visual Geometry Grounded Transformer (VGGT), a feed-forward network capable of reconstructing scenes from hundreds of views. VGGT outputs a full set of 3D attributes, including camera poses, depth maps, point maps, and 3D point tracks. While foundation models show great promise for 3D vision tasks, they are not yet specialized for visual localization, and have yet to show competitive performance with dedicated methods on large-scale benchmarks.
@@ 3 Problem statement
We first obtain the global descriptor $\mathbf{g}_{i}\in\mathbb{R}^{\text{n}}$ and local descriptors $\mathbf{l}_{ij}\in\mathbb{R}^{\text{m}}$ for each image $\mathit{I}_{i}$ and its $j$-th keypoint, respectively. The local descriptors $\mathbf{l}_{ij}$ are obtained using a pretrained off-the-shelf model [edstedt2024dedode]. The global descriptor $\mathbf{g}_{i}$ is a compact representation of the entire image, obtained via the aggregator module $f_{\text{agg}}$.
@@
where $K_{i}\in\mathbb{R}^{3\times 3}$ is the camera calibration matrix, $\mathbf{y}_{ij}\in\mathbb{R}^{2}$ denotes the $j$-th keypoint coordinate in $\mathit{I}_{i}$ and $\mathbf{x}_{ij}\in\mathbb{R}^{3}$ represents the $j$-th predicted scene coordinate. The hat operator $\hat{\cdot}$ denotes the homogeneous coordinate representation and $||\cdot||_{2}$ denotes the L2-norm. Although the optimization objective remains consistent with Eq. 1, more sophisticated objectives are typically used in practice to ensure stable training. For further details, we refer the reader to [brachmann2023accelerated, rscore].
Finally, the camera pose can be computed using pairs of 2D-3D correspondences $\{(\mathbf{y}_{ij},\mathbf{x}_{ij})\}$ using an off-the-shelf Perspective-n-Point solver [PoseLib, persson2018lambda].
@@ 4 Methodology
Figure 2: System overview. Our pipeline consists of four main components: (1) Local descriptor extraction using DeDoDe [edstedt2024dedode], (2) DINO [dino] feature extraction for visual representation, (3) Our proposed aggregator module that learns geometrically-consistent global descriptors from DINO features using dual consistency constraints, and (4) Scene coordinate regression model that predicts 3D coordinates from concatenated local-global descriptors.
Component 1: Local feature extraction. We extract sparse local descriptors using DeDoDe [edstedt2024dedode], which has demonstrated strong performance for coordinate regression tasks [rscore]. DeDoDe processes grayscale images to detect keypoints and outputs corresponding 256-dimensional descriptors. To reduce memory requirements and computational cost, we apply PCA [pca] compression to reduce descriptor dimensionality from 256 to 128 dimensions. Following [rscore], we employ the L-upright weights for keypoint detection and B-upright weights for descriptor extraction.
Component 2: Visual feature extraction. In parallel with local descriptor extraction, we extract dense visual features using a pretrained DINO encoder [dino, dinov2]. DINO features provide rich visual representations that capture semantic content across the entire image, complementing the sparse local information of local descriptors. The DINO encoder can be fine-tuned, though we keep them frozen in our experiments.
Component 3: Aggregator module. Our core contribution is the aggregator module that learns geometrically-consistent global descriptors from dense DINO features. Unlike existing approaches that derive global descriptors from geometrical cues alone, our aggregator enforces dual consistency by requiring both visual similarity and geometrical constraints from the covisibility graph before assigning similar global descriptors to image pairs. The aggregator follows similar architectural principles to SALAD [Izquierdo_CVPR_2024_SALAD] but incorporates our novel training strategy for geometrical consistency. The module first projects DINO descriptors into a lower-dimensional space and computes relevance scores for each descriptor with a scoring layer. These projected descriptors are then combined through a weighted aggregation to produce a compact descriptor $\mathbf{g}_{i}$. This descriptor is further compressed to $256$ dimensions using a PCA layer [pca]. The final product is concatenated with each local descriptor and passed to the fourth component, which predicts a 3D coordinate for each concatenated descriptor.
We use a modified version of the generalized contrastive loss [leyva2023gcl] for a given pair of global descriptors $\mathbf{g}_{i}$ and $\mathbf{g}_{j}$:
@@
Generally, batch composition is critical for training good global descriptors [izquierdo2024close, leyva2023gcl]. We divide our batch into three groups, each of size $b$ and sample $b$ positive pairs with overlap scores $\psi_{i,j}>0.5$, $b$ soft negative pairs with $0.25\leq\psi_{i,j}\leq 0.5$, and $b$ random pairs with $\psi_{i,j}=0$.
Component 4: Coordinate prediction. The final component performs scene coordinate regression using the fused local-global descriptors. We adopt the coordinate regression architecture from R-Score [rscore], which consists of multiple fully connected layers that map concatenated descriptors to 3D scene coordinates. These coordinates finally allow computing the camera pose $H_{i}$ for the input image $\mathit{I}_{i}$ using a Perspective-n-Point solver [PoseLib, persson2018lambda].
@@
AS (SIFT) [active_search] 750 MB 85.3 92.2 97.9 39.8 49.0 64.3 71.4
Cascaded [cheng2019cascaded] 140 MB 76.7 88.6 95.8 33.7 48.0 62.2 67.5
Squeezer [yang2022scenesqueezer] 240 MB 75.5 89.7 96.2 50.0 67.3 78.6 76.2
PixLoc [sarlin21pixloc] 2.13 GB 64.3 69.3 77.4 51.1 55.1 67.3 64.1
Learning-based methods ACE ($\times 50$) [brachmann2023accelerated] 205 MB 6.9 17.2 50.0 0.0 1.0 5.1 13.4
GLACE [GLACE2024CVPR] 27 MB 8.6 20.8 64.0 1.0 1.0 17.3 18.8
ESAC ($\times 50$) [brachmann2019expert] 1.31 GB 42.6 59.6 75.5 6.1 10.2 18.4 35.4
HSCNet++ [wang2024hscnet++] 274 MB 72.7 81.6 91.4 43.9 57.1 76.5 70.5
Neumap [tang2023neumap] 1.26 GB 80.8 90.9 95.6 48.0 67.3 87.8 78.4
@@
hLoc (D2-net) [hloc, dusmanu2019d2] 362 GB 78.0 82.8 88.0 84.2 89.8 92.0 73.7 79.3 87.2 83.9
DeViLoc [deviloc] $\approx$362 GB 86.9 91.5 96.3 88.7 93.7 96.1 78.5 84.2 93.7 90.0
ACE ($\times 50$) [brachmann2023accelerated] 205 MB 14.1 54.4 75.5 27.3 70.9 84.1 2.7 14.4 29.3 41.4
ESAC ($\times 50$) [brachmann2019expert] 1.4 GB 43.3 66.3 77.0 45.2 62.5 73.1 3.5 8.2 12.6 43.5
GLACE [GLACE2024CVPR] 42 MB 5.6 21.3 48.6 8.4 29.8 51.6 0.9 4.4 11.9 20.3
@@
R-Score (DeDoDe) [rscore, edstedt2024dedode] 102 MB 63.9 83.3 90.8 76.7 89.3 93.0 61.5 77.6 88.8 80.5

### v2

@@ Abstract
Figure 1: Performance overview on Aachen Day/Night dataset. Our method achieves significant improvements over existing learning-based approaches while maintaining comparable memory efficiency. Compared to R-Score [28], we achieve $6.1\%$ higher accuracy on night-time images at the $0.25\text{m}$ / $2^{\circ}$ threshold and 2.2% average improvement across all evaluation thresholds (detailed results in Table 1). This performance narrows the gap between learning-based and traditional structure-based methods while preserving the memory advantages of coordinate regression approaches.
@@ 1 Introduction
Visual localization, the task of determining a camera’s position and orientation from images alone, represents a core challenge in computer vision with critical applications in augmented reality, autonomous navigation, and robotics. While traditional structure-based approaches achieve robust performance through extensive feature matching and geometric verification, they impose significant computational and memory burdens that limit their practical deployment [48, 52].
Scene coordinate regression (SCR) methods [7, 60, 9, 41, 55, 59] have emerged as an attractive alternative, excelling in small-scale environments in both accuracy and memory efficiency. However, they face perceptual aliasing as a fundamental challenge in larger environments. In such environments, visually similar landmarks, such as repeated architectural elements, often generate nearly identical local descriptors. This similarity creates ambiguity that makes it difficult to associate descriptors with their correct 3D locations and severely hampers the training effectiveness.
To address the challenge of scene ambiguity, recent works [60, 28] have leveraged global context to disambiguate similar local features by incorporating global descriptors. R-Score [28] extracts geometric relationships from covisibility graphs via the node2vec optimizer [21]. In these graphs, edges represent spatial proximity between images, with images in close proximity in the covisibility graph sharing similar embeddings. These embeddings, when concatenated with local descriptors, enable effective disambiguation across scenes and substantially boost the performance of learning-based SCR methods.
However, R-Score derives global descriptors from geometric cues alone, decoupling them from visual content. This separation creates two fundamental problems: First, rich visual information that could enhance descriptor discriminability is ignored. Second, the system is vulnerable to the noise in the underlying covisibility graph, which is computed from potentially unreliable overlap scores [46]. When incorrect connections between dissimilar image pairs are formed, the resulting graph embeddings become partially misleading, degrading both descriptor quality and training stability.
@@
Our aggregator learns these geometrically-consistent global descriptors through a training strategy that combines an efficient batch mining strategy relying solely on the overlap scores, and thus eliminating the need for external place labels, with a modified Generalized Contrastive Loss [34] that uses a smoother distance function.
Experimental validation on challenging benchmarks demonstrates the effectiveness of our approach. As shown in Figure 1, our method achieves a $6.1\%$ improvement at the $0.25\text{m}$ / $2^{\circ}$ threshold compared to R-Score [28], with only a 6 MB increase in memory. This gain further narrows the gap to traditional structure-based methods while preserving the memory advantages that make SCR models attractive for practical deployment.
@@
    We demonstrate consistent improvements across challenging benchmarks: compared to R-Score [28], our method achieves average accuracy gains of $2.2\%$ and $2.5\%$ on the Aachen Day/Night [54] and the Hyundai Department Store datasets [32], respectively, while maintaining similar computational and memory efficiency.
@@ 2.1 Structure-based visual localization
Structure-based methods establish visual localization by matching query images against pre-built 3D scene representations. Early approaches [36, 52, 51] directly match 2D features with 3D point clouds to establish 2D-3D correspondences. These approaches typically construct a descriptor codebook for each 3D point by averaging descriptors from all database pixels in which the point appears. Direct matching methods can benefit from sophisticated search strategies based on both 2D-to-3D and 3D-to-2D search for additional matches [52, 51]. However, they remain vulnerable to perceptual aliasing, where visually similar but spatially distinct 3D points produce incorrect correspondences.
Hierarchical approaches [44, 48] address this limitation by leveraging global descriptors to streamline and refine the matching process. These approaches first retrieve database images similar to the query image using global descriptors, and then establish 2D-2D feature correspondences between retrieved and query images to obtain 2D-3D correspondences. This two-stage process significantly improves accuracy compared to direct matching.
Despite their strong performance, both direct and hierarchical approaches impose substantial memory requirements. They must maintain either complete database image descriptors [48, 44] or the descriptors for 3D points [36, 52, 51]. Although recent advances [62, 31] have explored descriptor compression, memory requirements of structure-based algorithms still scale linearly with environment size, unlike learning-based counterparts. This motivates our focus on learning-based alternatives that offer more favorable memory characteristics.
@@ 2.2 Learning-based visual localization
Scene coordinate regression (SCR) models represent a fundamentally different approach to visual localization, directly predicting 3D scene coordinates for pixels in query images [55, 59, 11, 13, 35]. Early SCR approaches employed random forests [55, 59], while subsequent approaches have adopted convolutional neural networks [11, 8, 9, 10, 16, 24] and fully-connected architectures [7, 41] to improve the prediction accuracy and computational efficiency.
However, SCR approaches have historically underperformed in large-scale environments due to perceptual aliasing and ambiguous visual patterns. This fundamental challenge has motivated recent efforts to incorporate global context into SCR architectures. Recent works [60, 28] integrate global descriptors to provide contextual information for disambiguation. GLACE [60] incorporates off-the-shelf descriptors with noise augmentation during training, but achieves limited robustness in large scenes due to insufficient consistency between descriptors of the same location. R-Score [28] improves scalability by learning global descriptors from a covisibility graph using node2vec embeddings [21]. While R-Score demonstrated significant improvements, its graph-based representation operates independently of visual content in the image space, potentially missing important visual cues that could enhance descriptor discriminability.
@@ 2.3 Local descriptors
Local feature descriptors form the foundation of many visual localization systems, enabling the identification and description of consistent pixels under variations in lighting, viewpoint, and scale. Classical methods [37, 4] detect invariant keypoints that can be reliably matched across different views. These approaches achieve strong real-world performance with computational efficiency, making them popular choices for structure-based localization systems [36, 52, 25, 53].
More recent work has shifted toward learning-based local descriptors [15, 42, 18, 19]. Noh et al. [42] introduces an attention mechanism to highlight semantically meaningful local features while also estimating their confidence. SuperPoint [15] learns keypoint detection and description on a synthetic dataset containing simple geometric shapes. D2-Net [18] employs a unified convolutional neural network for both dense feature description and keypoint detection, postponing the detection process to produce more stable keypoints compared to traditional methods that rely on early detection of low-level image structures. R2D2 [47] jointly learns keypoint detection, descriptor extraction, and discriminativeness prediction. This integrated approach reduces the impact of ambiguous regions, leading to more robust and reliable keypoints and descriptors. DeDoDe [19] trains a detector using tracks from large-scale structure-from-motion and learns descriptors by optimizing a mutual nearest neighbor objective over keypoints. DeDoDe achieves strong performance in feature matching and has proven particularly effective for scene coordinate regression models [28].
@@ 2.4 Global descriptors
Global descriptors enable finding the most similar database images to an input query image by encoding image-wide visual characteristics into compact representations [39, 38]. Current systems often reduce this problem to a similarity search in a $d$-dimensional descriptor space [23, 3, 1, 6, 40, 58, 27, 26, 34].
Global descriptors can be obtained by aggregating either local descriptors [40, 58], multiple convolutional neural network layers [3, 1, 6], or DINO features [12, 43] via optimal transport [27], into a single global descriptor vector. In the context of scene coordinate regression, global descriptors have emerged as crucial components providing contextual information for disambiguating local descriptors in the presence of perceptual aliasing [60, 28]. Our work advances this trend by introducing a dedicated aggregator module that produces high-quality global descriptors through joint optimization with both geometric and visual constraints, leading to further improvements in scene coordinate regression performance.
@@ 2.5 Foundation models
Foundation models [64, 61, 33, 17] have demonstrated strong capabilities in estimating various 3D properties, such as point clouds, depth maps, and camera poses, from just a few images, thanks to their high-quality representations. Dust3r [64] pioneered this line of work by introducing pointmap regression using powerful pretrained features [65]. Building on this trend, Wang et al. [61] proposed the Visual Geometry Grounded Transformer (VGGT), a feed-forward network capable of reconstructing scenes from hundreds of views. VGGT outputs a full set of 3D attributes, including camera poses, depth maps, point maps, and 3D point tracks. While foundation models show great promise for 3D vision tasks, they are not yet specialized for visual localization, and have yet to show competitive performance with dedicated methods on large-scale benchmarks.
@@ 3 Problem statement
We first obtain the global descriptor $\mathbf{g}_{i}\in\mathbb{R}^{\text{n}}$ and local descriptors $\mathbf{l}_{ij}\in\mathbb{R}^{\text{m}}$ for each image $\mathit{I}_{i}$ and its $j$-th keypoint, respectively. The local descriptors $\mathbf{l}_{ij}$ are obtained using a pretrained off-the-shelf model [19]. The global descriptor $\mathbf{g}_{i}$ is a compact representation of the entire image, obtained via the aggregator module $f_{\text{agg}}$.
@@
where $K_{i}\in\mathbb{R}^{3\times 3}$ is the camera calibration matrix, $\mathbf{y}_{ij}\in\mathbb{R}^{2}$ denotes the $j$-th keypoint coordinate in $\mathit{I}_{i}$ and $\mathbf{x}_{ij}\in\mathbb{R}^{3}$ represents the $j$-th predicted scene coordinate. The hat operator $\hat{\cdot}$ denotes the homogeneous coordinate representation and $||\cdot||_{2}$ denotes the L2-norm. Although the optimization objective remains consistent with Eq. 1, more sophisticated objectives are typically used in practice to ensure stable training. For further details, we refer the reader to [7, 28].
Finally, the camera pose can be computed using pairs of 2D-3D correspondences $\{(\mathbf{y}_{ij},\mathbf{x}_{ij})\}$ using an off-the-shelf Perspective-n-Point solver [30, 45].
@@ 4 Methodology
Figure 2: System overview. Our pipeline consists of four main components: (1) Local descriptor extraction using DeDoDe [19], (2) DINO [12] feature extraction for visual representation, (3) Our proposed aggregator module that learns geometrically-consistent global descriptors from DINO features using dual consistency constraints, and (4) Scene coordinate regression model that predicts 3D coordinates from concatenated local-global descriptors.
Component 1: Local feature extraction. We extract sparse local descriptors using DeDoDe [19], which has demonstrated strong performance for coordinate regression tasks [28]. DeDoDe processes grayscale images to detect keypoints and outputs corresponding 256-dimensional descriptors. To reduce memory requirements and computational cost, we apply PCA [22] compression to reduce descriptor dimensionality from 256 to 128 dimensions. Following [28], we employ the L-upright weights for keypoint detection and B-upright weights for descriptor extraction.
Component 2: Visual feature extraction. In parallel with local descriptor extraction, we extract dense visual features using a pretrained DINO encoder [12, 43]. DINO features provide rich visual representations that capture semantic content across the entire image, complementing the sparse local information of local descriptors. The DINO encoder can be fine-tuned, though we keep them frozen in our experiments.
Component 3: Aggregator module. Our core contribution is the aggregator module that learns geometrically-consistent global descriptors from dense DINO features. Unlike existing approaches that derive global descriptors from geometrical cues alone, our aggregator enforces dual consistency by requiring both visual similarity and geometrical constraints from the covisibility graph before assigning similar global descriptors to image pairs. The aggregator follows similar architectural principles to SALAD [27] but incorporates our novel training strategy for geometrical consistency. The module first projects DINO descriptors into a lower-dimensional space and computes relevance scores for each descriptor with a scoring layer. These projected descriptors are then combined through a weighted aggregation to produce a compact descriptor $\mathbf{g}_{i}$. This descriptor is further compressed to $256$ dimensions using a PCA layer [22]. The final product is concatenated with each local descriptor and passed to the fourth component, which predicts a 3D coordinate for each concatenated descriptor.
We use a modified version of the generalized contrastive loss [34] for a given pair of global descriptors $\mathbf{g}_{i}$ and $\mathbf{g}_{j}$:
@@
Generally, batch composition is critical for training good global descriptors [26, 34]. We divide our batch into three groups, each of size $b$ and sample $b$ positive pairs with overlap scores $\psi_{i,j}>0.5$, $b$ soft negative pairs with $0.25\leq\psi_{i,j}\leq 0.5$, and $b$ random pairs with $\psi_{i,j}=0$.
Component 4: Coordinate prediction. The final component performs scene coordinate regression using the fused local-global descriptors. We adopt the coordinate regression architecture from R-Score [28], which consists of multiple fully connected layers that map concatenated descriptors to 3D scene coordinates. These coordinates finally allow computing the camera pose $H_{i}$ for the input image $\mathit{I}_{i}$ using a Perspective-n-Point solver [30, 45].
@@
AS (SIFT) [52] 750 MB 85.3 92.2 97.9 39.8 49.0 64.3 71.4
Cascaded [14] 140 MB 76.7 88.6 95.8 33.7 48.0 62.2 67.5
Squeezer [66] 240 MB 75.5 89.7 96.2 50.0 67.3 78.6 76.2
PixLoc [50] 2.13 GB 64.3 69.3 77.4 51.1 55.1 67.3 64.1
Learning-based methods ACE ($\times 50$) [7] 205 MB 6.9 17.2 50.0 0.0 1.0 5.1 13.4
GLACE [60] 27 MB 8.6 20.8 64.0 1.0 1.0 17.3 18.8
ESAC ($\times 50$) [10] 1.31 GB 42.6 59.6 75.5 6.1 10.2 18.4 35.4
HSCNet++ [63] 274 MB 72.7 81.6 91.4 43.9 57.1 76.5 70.5
Neumap [57] 1.26 GB 80.8 90.9 95.6 48.0 67.3 87.8 78.4
@@
hLoc (D2-net) [48, 18] 362 GB 78.0 82.8 88.0 84.2 89.8 92.0 73.7 79.3 87.2 83.9
DeViLoc [20] $\approx$362 GB 86.9 91.5 96.3 88.7 93.7 96.1 78.5 84.2 93.7 90.0
ACE ($\times 50$) [7] 205 MB 14.1 54.4 75.5 27.3 70.9 84.1 2.7 14.4 29.3 41.4
ESAC ($\times 50$) [10] 1.4 GB 43.3 66.3 77.0 45.2 62.5 73.1 3.5 8.2 12.6 43.5
GLACE [60] 42 MB 5.6 21.3 48.6 8.4 29.8 51.6 0.9 4.4 11.9 20.3
@@
R-Score (DeDoDe) [28, 19] 102 MB 63.9 83.3 90.8 76.7 89.3 93.0 61.5 77.6 88.8 80.5